# Probabilistic Robustness Analysis in High Dimensional Space: Application to Semantic Segmentation Network

### v1

@@ 4 Experiments
Table 1: Model details and probabilistic guarantees. Given hyperparameters $\ell$, $m$, and coverage $\delta_{1}=1-\epsilon$, the confidence $\delta_{2}$ is computed for each experiment. The reported verification runtime is averaged over all experiments run on a model. One important point is, our verification runtime depends mainly on inference time, and hyperparameters ($\ell,m$), and is only slightly sensitive to perturbation magnitude and dimension $(e,r)$. For per-experiment runtimes, see Figure 5 in Appendix E.
@@
Cityscapes (Appendix E) HRNetV2 $1024\!\times\!2048\!\times\!3$ $256\!\times\!512\!\times\!19$ $65,859,379$ $(921,1)$ 0.99 0.997 3.5 min (Naive)
RQ1: How does the technique scale with model size, number of perturbed pixels, and perturbation level? How does output high dimensionality, and the irregularity in output distribution of large and complex NNs make our guarantees overly conservative?
To assess scalability, we test our verification method on high-dimensional datasets (e.g., Lung Segmentation (Jaeger et al., 2014; Candemir et al., 2014), OCTA-500 (Li et al., 2019), CamVid (Brostow et al., 2009)) using large pre-trained models. We evaluate performance on a subset of $200$ test images across varying perturbation dimensions and magnitudes, with results shown in Figure 2. Model details and probabilistic guarantees are summarized in Table 1, where we show the following $\langle\epsilon,\ell,m\rangle$ guarantee:
@@
In these experiments, we studied an $r$-dimensional darkening adversary (see Figure 1), where $r^{\prime}$ pixels in an image $x$ with intensity above $150/255$ in all channels are randomly selected for perturbation ($r=nc\times r^{\prime}$). Each direction $x_{i}^{\text{noise}}$ corresponds to darkening a channel of one such pixels, and $x^{\text{adv}}$ is parameterized by an $r$-dimensional coefficient vector $\lambda\in[\underline{\lambda},\overline{\lambda}]$. Here, $\overline{\lambda}$ induces full darkening (intensity zero), while $\underline{\lambda}$ applies partial darkening. This bound defines the perturbation space $x^{\text{adv}}\in\mathbf{I}$. We also present a demo for the status of all pixels in the segmentation mask in Figures 6,7 and 8.
@@
The degree of conservatism can be assessed by comparing $\underline{\hat{y}},\overline{\hat{y}}$] with $\underline{y},\overline{y}]$. To measure this, we compute $\mathsf{bound\_ratio}=\sum_{k=1}^{h\times w\times L}(\overline{\hat{y}}(k)-\underline{\hat{y}}(k))/\sum_{k=1}^{h\times w\times L}(\overline{y}(k)-\underline{y}(k))$ and due to the high cost of $10^{6}$ inferences on the models, we only perform the conservatism analysis on one specific case with BiSeNet from Figure 2, Table 2 in Appendix E details the numerical results.
@@ 5 Conclusion
Our probabilistic verification method is both scalable and data-efficient. While it does not offer deterministic guarantees, it can provide meaningful assurances in regions where existing deterministic tools are not applicable due to computational challenges. In this paper, we also demonstrated this capability through a series of numerical experiments.
Appendix A Double-Step Probabilistic Guarantee
The necessary steps to establish the two-sided probabilistic guarantee, along with the flexibility in choosing the dataset size $m$, rank $\ell$, and failure probability $\epsilon$, begin with the theory of order statistics David and Nagaraja (2004). This flexibility means that the double-step probability constraint remains valid regardless of the particular values chosen for $m$, $\ell\leq m$, and $\epsilon\in[0,1]$. The following lemma introduces the concept of order statistics.
Lemma 1 (From David and Nagaraja (2004))
Consider $m$ independent and identically distributed (i.i.d.), real-valued data points $x\in[0,1]$ drawn from uniform distribution $x\sim\mathcal{U}$. Suppose we sort them in ascending order and denote the $\ell^{th}$ smallest number by $x_{\ell}$,(i.e., we have $x_{1}<x_{2}<\ldots<x_{m}$). Let $\mathbf{Beta}(\alpha,\beta)$ denote the Beta distribution [^3] The Beta distribution is a family of continuous probability distributions defined on the interval $0\leq x\leq 1$ with shape parameters $\alpha$ and $\beta$, and with probability density function $f(x;\alpha,\beta)=\frac{x^{\alpha-1}(1-x)^{\beta-1}}{B(\alpha,\beta)}$, where the constant $B(\alpha,\beta)=\frac{\Gamma(\alpha)\Gamma(\beta)}{\Gamma(\alpha+\beta)}$ and $\Gamma(z)=\int_{0}^{\infty}t^{z-1}e^{-t}\mathbf{d}t$ is the Gamma function.. Then the uniform random variable at rank $\ell$ follows:
$x_{\ell}\sim\mathbf{Beta}(\ell,m+1-\ell)$
In the context of conformal inference Lemma 1, has been extended from a uniform and domain-bounded distribution to a general distribution as follows:
Lemma 2 (From Vovk et al. (2005))
Consider $m$ independent and identically distributed (i.i.d.), real-valued data points $R$ drawn from some distribution $R\sim\mathcal{D}$. Suppose we sort them in ascending order and denote the $\ell^{th}$ smallest number by $R_{\ell}$, (i.e., we have $R_{1}<R_{2}<\ldots<R_{m}$). Let $\mathbf{Beta}(\alpha,\beta)$ denote the Beta distribution. Then, for an arbitrary $R_{m+1}$ drawn from the same distribution $\mathcal{D}$, the following holds:
$\Pr\left[R_{m+1}<R_{\ell}\right]\sim\mathbf{Beta}(\ell,m+1-\ell),\quad 1\leq\ell\leq m.$ (11)
This result is motivated by two key observations: (1) for a continuous distribution $R\sim\mathcal{D}$, the cumulative probability $\Pr[R\leq r]$ is uniformly distributed on $[0,1]$, and (2) sorting the samples $R_{1}<R_{2}<\ldots<R_{m}$ implies that the corresponding cumulative probabilities $\Pr[R\leq R_{\ell}]$ are also ordered, i.e., $\Pr[R\leq R_{1}]<\Pr[R\leq R_{2}]<\ldots<\Pr[R\leq R_{m}]$. These observations implies that for all $\ell=0,1,2,\ldots,m$, the random variable $\delta:=\Pr[R\leq R_{\ell}]$ satisfies the conditions outlined in Lemma 1.
In simple terms, this means that while a fixed dataset $R_{1},R_{2},\ldots,R_{m}$ is not sufficient to determine the exact value of $\delta=\Pr[R<R_{\ell}]$, we can infer that the value of $\delta$ must follow a Beta distribution that is solely defined based on $m$ and $\ell$. This allows us to tune the parameters $m$ and $\ell$ to provide strong guarantees for the coverage $R<R_{\ell}$ where $R_{\ell}$ is obtained from our fixed sampled dataset.
This guarantee can be equivalently expressed as a two-sided probabilistic bound. Specifically, let $\mathsf{betacdf}_{\delta}(\ell,m+1-\ell)$ denote the cumulative distribution function (CDF) of the Beta distribution—also known as the regularized incomplete beta function—evaluated at $\delta$. Then, based on the definition of cumulative density function, CDF, for any confidence $\delta_{1}\in[0,1]$ and any arbitrary distribution $R\sim\mathcal{D}$, the following inequality holds:
$\Pr\left[\ \ \Pr[\ \ R\leq R_{\ell}\ \ ]>\delta_{1}\ \ \right]>1-\mathsf{betacdf}_{\delta_{1}}(\ell,m+1-\ell).$
Note that the key assumption is that all variables $R,R_{1},R_{2},\ldots,R_{m}$ are independently drawn from the same distribution $\mathcal{D}$, and that the previously sampled values are sorted such that $R_{1}<R_{2}<\ldots<R_{m}$. Importantly, the resulting probabilistic guarantee remains valid for any arbitrary choice of sample size $m$, rank $\ell$, and confidence level $\delta_{1}$, provided they lie within appropriate bounds.
Appendix B A Clarifying Example for Conformal Inference
Assume an arbitrary distribution $R\sim\mathcal{D}$. We sample a calibration dataset of size $m=100,000$, we sort the samples as $R_{1}<R_{2}<\ldots<R_{100,000}$ and we select the sample at rank $\ell=99,999$. In this case, given an unseen draw $R^{\text{unseen}}$ from $\mathcal{D}$ for a miscoverage level $\epsilon=0.0001$, the $\langle\epsilon,\ell,m\rangle$ guarantee for logical statement $P(\ell,m)=[R^{\text{unseen}}<R_{\ell}]$ is:
$\Pr\left[\ \ \Pr[\ \ R^{\text{unseen}}<R_{99,999}\ \ ]\ >\ 0.9999\ \ \right]>\ 0.9995008$
which is valid for any calibration dataset $\mathbf{M}=\left\{R_{1},R_{2},\ldots,R_{100,000}\right\}$ sampled from distribution $\mathcal{D}$ if and only if $R^{\text{unseen}}$ is also sampled from $\mathcal{D}$, and $R_{99,999}$ is the $99,999^{\text{th}}$ smallest member of $\mathbf{M}$.
Appendix C A Clarifying Example on Naive Approach
Figure 3: The figure depicts two reachable sets corresponding to different $\langle\epsilon,\ell,m\rangle$ guarantees. Both $S_{1}$ and $S_{2}$ are obtained using the naïve approach, with $S_{2}$ providing a stronger guarantee than $S_{1}$.
Since the naive technique is architecture-agnostic we apply that for the reachability analysis of a feedforward $\mathrm{ReLU}$ neural network with $60$ hidden-layers of width $100$, an input layer of dimension $784$ and an output layer of dimension $2$. The set of inputs is $\mathbf{I}:=[0,1]^{784}$ and we generate both calibration and train datasets by sampling $\mathbf{I}$ uniformly. e.g., $\mathcal{W},\mathcal{W}^{\prime}$ are both uniform distributions.
1:  We firstly generate a calibration dataset of size $m=200,000$ and a train dataset of size $t=10,000$, we also consider the rank $\ell=199,998$ and we target the miscoverage level of $\epsilon=0.0001$. In this case we compute a reachable set, $\mathbf{S}_{1}$ with the following $\langle\epsilon,\ell,m\rangle$ guarantee.
2:  Secondly, we set $m=8464287$, $\ell=8464286$ and $\epsilon=0.000002$ which results in another reachset $\mathbf{S}_{2}$ with the following $\langle\epsilon,\ell,m\rangle$ guarantee.
$x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}\ \ \Rightarrow\ \ \Pr\left[\ \ \Pr[\ \ f(x)\in\mathbf{S}_{2}\ \ ]>0.999998\ \ \right]>0.9999992.$ (13)
Figure 4 shows the reachable sets $\mathbf{S}_{1}$ and $\mathbf{S}_{2}$ in the presence of $10^{6}$ new simulations, $f(x),x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$, for validation. Our calculations show that the proportion of points lying outside of $\mathbf{S}_{1}$ and $\mathbf{S}_{2}$ are 0.000013 and 0.000001, respectively, which aligns well with the proposed $\langle\epsilon,\ell,m\rangle$ guarantees.
Remark 1
One might argue that the dimensionality issue can be addressed by partitioning the output layer and performing reachability analysis on each partition separately. However, this approach is suboptimal, as the overall probabilistic guarantee deteriorates. Specifically, the combined guarantee is governed by a union bound over the individual partitions, which can lead to an unacceptably loose $\langle\epsilon,\ell,m\rangle$ guarantee.
Appendix D More Details on Clipping Block, Surrogate Model & Minkowski Sum
$\mathbf{CH}=\sum_{j=1}^{t}\alpha_{j}\mathbf{vec}(y_{j}^{\text{train}}),\qquad\mathbf{s.t.}\quad\sum_{j=1}^{t}\alpha_{j}=1,\quad\alpha_{1},\cdots,\alpha_{t}>0$
that for $l=2$ is a quadratic programming and for $l=1,\infty$ is a linear programming. Among the three options, we are most interested in $l=\infty$, since it minimizes the maximum error across the logit components, making it a suitable choice for reachability analysis. In addition, the $l=2$ formulation is far more memory-intensive compared to the $l=\infty$ case. When we ran it on the OCTA-500 dataset using the Vortex machine (540 GB RAM), it resulted in an out-of-memory error. We also attempted the PGD variant of quadratic programming for $l=2$; however, this led to severe CPU thread oversubscription issues. Thus, the LP formulation scales more gracefully. Each problem is linear, and batching helps amortize model builds, and parallelism is also effective — which is why we observe it running stably on our setup.
Thus, our surrogate model $g$ is composed of two main components. The first is the base model $f$, which takes the vectorized image $\mathbf{vec}(x^{\text{unseen}})$ and produces the corresponding logit vector $\mathbf{vec}(y^{\text{unseen}})$. The second is the clipping block, which takes $\mathbf{vec}(y^{\text{unseen}})$ as input and returns its projection onto the convex hull $\mathbf{CH}$, denoted as $\mathbf{vec}(\hat{y}^{\text{unseen}})$. In other words,
$g(\mathbf{vec}(x^{\text{unseen}}))=\mathbf{vec}(\hat{y}^{\text{unseen}})$
and obviously $\mathbf{CH}$ represents the deterministic reachable set of the surrogate model, $\mathbf{R}_{g}(\mathbf{I})=\mathbf{CH}.$
To generate the convex hull, it is not necessary to include the entire training dataset. Instead, one can identify the extreme points Sartipizadeh and Vincent (2016) and construct the convex hull using only those points, which reduces the number of coefficients $\alpha$ without sacrificing accuracy. This approach makes the convex hull more efficient to be used as it reduces the number of coefficients $\alpha$. Existing algorithms for identifying extreme points include Quickhull Barber et al. (1996) and Clarkson’s algorithm Clarkson (1988). Their computational complexities are $\mathcal{O}(t\log(t))$ for $n=2,3$ and $\mathcal{O}(t^{\lfloor n/2\rfloor})$ for $n\geq 4$, which do not scale well to high-dimensional spaces. Another approach to this problem is to approximate the convex hull by allowing a controlled level of inaccuracy Sartipizadeh and Vincent (2016); Jia et al. (2022). However, as reported in the literature Casadio et al. (2025), this method has not shown promising performance even in high-dimensional settings such as large language models, whose dimensionality is still lower than that of semantic segmentation networks.
For this reason, we keep the original formulation and retain all points in the training dataset when constructing the convex hull. Importantly, this does not create scalability issues for our experiments, since conformal inference is highly data-efficient. For example, guarantees of order $\delta_{1}=0.999$ and $\delta_{2}=0.997$ can be verified with a calibration dataset of size $m=8000$. Because we consider the size of the training dataset to be at most $50\%$ of the calibration dataset size, then $t=4000$ training samples suffice for our experiments, requiring only $4000$ coefficients $\alpha$ in the convex hull formulation. Moreover, since the number of extreme points grows exponentially with the dimension of the logits, i.e., $\mathcal{O}((\log t)^{n-1})$ Dwyer (1988), it is highly likely that most training samples will serve as extreme points. This renders attempts to distinguish between extreme and interior points largely ineffective.
[…]
$\mathbf{HR}=\mathcal{R}_{q}^{\epsilon}(\mathcal{W};\ \ell,m)=c+\sum_{i=1}^{n}\beta_{i}\sigma_{i}e_{i},\qquad\mathbf{s.t.}\quad\sum_{i=1}^{n}\beta_{i}=1,\quad\beta_{1},\cdots,\beta_{n}>0$
$\mathcal{R}_{f}^{\epsilon}(\mathcal{W};\ell,m)=\mathcal{R}_{g}(\mathbf{I})\oplus\mathcal{R}_{q}^{\epsilon}(\mathcal{W};\ \ell,m)=\mathbf{CH}\oplus\mathbf{HR}=c+\sum_{j=1}^{t}\alpha_{j}A\ v^{\text{train}}_{j}+\sum_{i=1}^{n}\beta_{i}\sigma_{i}e_{i},$
$\text{where,}\qquad\alpha_{1},\cdots,\alpha_{t},\beta_{1},\cdots,\beta_{n}\geq 0,\quad\sum_{j=1}^{t}\alpha_{j}=1,\quad\sum_{i=1}^{n}\beta_{i}=1.$
The projection of the probabilistic reach set $\mathcal{R}_{f}^{\epsilon}(\mathcal{W};\ell,m)$ onto each class of a given pixel can be performed efficiently. To this end, we collect the training points $v_{j}^{\text{train}},;j=1,\ldots,t$. Since the mapping from latent space to the original space is linear ($\mathbf{vec}(\hat{y})=Av$), the vertices of the convex hull in latent space are preserved under this transformation. In other words, if $v^{*}$ is an extreme point of the convex hull in latent space, then $Av^{*}$ is an extreme point of the convex hull in the original space. Therefore, we collect the points $\mathbf{vec}(\hat{y}_{j}^{\text{train}})=Av_{j}^{\text{train}}$ and compare them to determine the upper bound $\mathbf{ub}$ and lower bound $\mathbf{lb}$ of the surrogate model’s predictions. Consequently, the clipping block in the surrogate model ensures that every unseen point $\mathbf{vec}(\hat{y}^{\text{unseen}})$ satisfies
$\mathbf{lb}\leq\mathbf{vec}(\hat{y}^{\text{unseen}})<\mathbf{ub}.$
For a given pixel and class, let the corresponding logit value be the $k$-th component of $\mathbf{vec}(y^{\text{unseen}})$. Then the projection of the reachset on this component satisfies
$c(k)+\mathbf{lb}(k)-\sigma_{k}\leq\mathbf{vec}(y^{\text{unseen}})(k)\leq c(k)+\mathbf{ub}(k)+\sigma_{k}.$
We conclude this section with a comparison against our naïve approach. Specifically, we consider the experiment described in Appendix C for the given guarantee in Eq. (12), using a calibration dataset of size m=200,000m=200{,}000 and compare the resulting bounds. For this comparison, we employ a clipping block that projects with $l=\infty$. We constructed the convex hull using a training dataset of size $t=4000$. In addition, we sampled a separate dataset of size $t^{\prime}=10000$ to compute the normalization factors (see Eq. (8)) prior to performing conformal inference for obtaining the inflating hyper-rectangle. The resulting reachable set, together with the outcome of the naïve approach, in the presence of $10^{6}$ new simulations, $f(x),x\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{I}$, are shown in Figure 4. Our calculation shows the proportion of the points lying outside of the reachset is only $0.000012$, which aligns well with the proposed $\langle\epsilon,\ell,m\rangle$ guarantee in Eq. (12).
Appendix E Additional Research Questions
Table 2: This table reports the bound ratio, $\mathsf{bound\_ratio}$, and the empirical miscoverage level of our reachset, $\hat{\epsilon}$ for a selection of test images, perturbation level and perturbation dimensions. We performed $10^{6}$ inferences to estimate the conservatism. Observing a nonzero $\hat{\epsilon}$ is a promising indication of a low level of conservatism.
(a) Image, 0001TP_008790.png
(b) Predicted Mask from BiSeNet
(c) Image perturbed over $6912$ darkened pixels.
(d) Image perturbed over $41472$ darkened pixels.
Figure 6: Visualization for verification on BiSeNet for a test image from the CamVid dataset. (a) Test image. (b) Segmentation mask predicted by BiSeNet. (c,d) Pixels selected for perturbation (in black) on the test image (we sampled $6192$ (1%) and $41472$ (6%) pixels with R, G, B intensities above $150/255$). (e,f) Robust (white), non-robust (red), and unknown (black) pixels for perturbation magnitudes $e=5/255$ and dimensions $r=6192\times 3$ and $r=41472\times 3$.
(a) Image, CHNCXR_0005_0.png
(b) Predicted Mask from UNet1
(c) Image perturbed over $2621$ darkened pixels.
(d) Image perturbed over $15726$ darkened pixels.
Figure 7: Visualization for verification on UNet1 for a test image from the Lung Segmentation dataset. (a) Test image. (b) Segmentation mask predicted by UNet1. (c,d) Pixels selected for perturbation (in black) on the test image (we sampled $2621$ (1%) and $15726$ (6%) pixels with Gray intensities above $150/255$). (e,f) Robust (white), non-robust (red), and unknown (black) pixels for perturbation magnitudes $e=5/255$ and dimensions $r=2621\times 1$ and $r=15726\times 1$.
(a) Image, 10491.bmp from OCTA-500 dataset.
(b) Predicted Mask from UNet2
(c) Image perturbed over $924$ darkened pixels.
(d) Image perturbed over $5544$ darkened pixels.
Figure 8: Visualization for verification on UNet2 for a test image from the OCTA-500 dataset. (a) Test image. (b) Segmentation mask predicted by UNet2. (c,d) Pixels selected for perturbation (in black) on the test image (we sampled $924$ (1%) and $5544$ (6%) pixels with Gray intensities above $150/255$). (e,f) Robust (white), non-robust (red), and unknown (black) pixels for perturbation magnitudes $e=5/255$ and dimensions $r=924\times 1$ and $r=5544\times 1$.
We applied techniques such as $\alpha$-$\beta$-CROWN (Zhou et al., 2024) and NNV (Tran et al., 2021) to our segmentation experiments using the UNet1, UNet2, and BiSeNet models in RQ1. However, due to the size of these models and the high perturbation levels considered, both methods encountered out-of-memory errors and failed to complete verification on the same hardware used for our approach. This highlights that while deterministic guarantees are often preferable, they may not always be computationally practical. In such cases, our probabilistic verification method for SSNs offers a scalable and effective alternative.
RQ3: How Does the Performance of our Methodology Compare to Existing Probabilistic Verification Methods for Neural Networks in the Literature?
In this section, we present a comparison with the works of Fiacchini and Alamo (2021); Anani et al. (2024) on the Cityscapes dataset Cordts et al. (2016). Since the formulation of verification guarantees differs slightly across these approaches, we first provide a general overview of the techniques proposed in Fiacchini and Alamo (2021); Anani et al. (2024), followed by a brief description of our own method. This will clarify how the comparison can be meaningfully established.
Overview 1. In Fischer et al. (2021), the authors introduced a probabilistic method to verify Semantic Segmentation Neural Networks. Inspired by Cohen et al. (2019), they introduced randomness via a Gaussian noise $\nu\sim\mathcal{N}(0,\sigma^{2})\in\mathbb{R}^{h\times w\times nc}$, applied to the input, and constructed a smoothed version of the segmentation model, denoted $\overline{\mathsf{SSN}}(x)(i,j)$, for each mask pixel $(i,j)$:
@@ Appendix E Additional Research Questions
This leads to a unified radius $r=\sigma\Phi^{-1}(\tau)$ for the input space. While this simplifies the formulation, it introduces the concept of abstained or uncertifiable pixels—those that cannot meet the desired confidence threshold. To mitigate the issue of multiple comparisons (i.e., union bounds), the authors propose statistical corrections such as the Bonferroni and Holm–Bonferroni methods CE (1936); Holm (1979). Assuming the set of certifiable pixels is defined as $\mathbf{CERT}=\left\{(i,j)\mid(i,j)\text{ is certifiable}\right\}$, they propose the final guarantee in the following form:
@@
This method is referred to as $\mathsf{SEGCERTIFY}$. A key limitation of this technique is the presence of abstained pixels. The authors in Anani et al. (2024) addressed this by proposing an adaptive approach, $\mathsf{ADAPTIVECERTIFY}$, which reduces the number of uncertifiable pixels. However, the guarantees proposed in Fiacchini and Alamo (2021); Anani et al. (2024) apply only to the smoothed model, not the base model. To express this in terms of the base model, we define the following problem:
@@
Overview 2. In our approach, for a given image $x$ and input set $\mathbf{B}_{r}(x)$, we perform a probabilistic reachability analysis with a $\langle\epsilon,\ell,m\rangle$ guarantee. Due to the presence of conservatism in our reachability technique, some mask pixels are marked as robust (i.e., certifiable), while others cover multiple classes and are considered unknown or uncertifiable. Let $\mathbf{CONFORMAL\_CERT}=\left\{(i,j)\mid(i,j)\text{ is certifiable}\right\}$. Then our verification objective is:
Problem 2. Given an image $x$, input set $\mathbf{B}_{r}(x)$, and sampling distribution $x^{\prime}\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{B}_{r}(x)$, for coverage level $\delta_{1}=1-\epsilon\in[0,1]$, hyper-parameters $\ell,m$ and confidence level $\delta_{2}=1-\mathsf{betacdf}_{\delta_{1}}(\ell,m+1-\ell)$, we aim to show the following $\langle\epsilon,\ell,m\rangle$ guarantee, where here the set $\mathbf{CONFORMAL\_CERT}$ will be also determined through the verification process:
@@
Comparison. The unknown pixels in our method are conceptually equivalent to the abstained pixels in Fiacchini and Alamo (2021); Anani et al. (2024). Thus, given the same input set $\mathbf{B}_{r}(x)$ where $r$ is provided by Fiacchini and Alamo (2021), the comparison focuses to show which technique results in fewer uncertifiable mask pixels. To this end, we replicate the setup of Table 1 from Anani et al. (2024), using the same HrNetV2 model Wang et al. (2020)trained on the Cityscapes dataset with the HrNetV2-W48 backbone. In this case study, the results of our Naive approach were approximately equivalent to those of the surrogate-based approach. Therefore, we use the Naive technique for comparison, as it is more efficient than the surrogate-based method. We compare the number of uncertifiable pixels produced by our naive approach against the abstained pixels reported in their results. Our findings are summarized in Table 3, which shows the average over 200 test images. We also report our verification runtime for completeness.
Table 3: Percentage of uncertifiable pixels under different $(\sigma,\tau,r)$ settings for Fiacchini and Alamo (2021), Anani et al. (2024), and our Naive approach. The verification runtime is $210$ seconds for all experiments.
$\sigma$ $\tau$ $r$ Fiacchini and Alamo (2021) (%) Anani et al. (2024) (%) Ours (%)

### v2

@@ 4 Experiments
Table 1: Model details and probabilistic guarantees. Given hyperparameters $\ell$, $m$, and coverage $\delta_{1}=1-\epsilon$, the confidence $\delta_{2}$ is computed for each experiment. The reported verification runtime is averaged over all experiments run on a model. One important point is, our verification runtime depends mainly on inference time, and hyperparameters ($\ell,m$), and is only slightly sensitive to perturbation magnitude and dimension $(e,r)$. For per-experiment runtimes, see Figure 4 in Appendix D.
@@
Cityscapes (Appendix D) HRNetV2 $1024\!\times\!2048\!\times\!3$ $256\!\times\!512\!\times\!19$ $65,859,379$ $(921,1)$ 0.99 0.997 3.5 min (Naive)
RQ1: How does the technique scale with model size and dimensionality, number of perturbed pixels, and perturbation level?
To assess scalability, we test our verification method on high-dimensional datasets (e.g., Lung Segmentation (Jaeger2014lungdata1; Candemir2014lungdata2), OCTA-500 (j.media.2024.103092-19), CamVid (Brostow2009camvid)) using large pre-trained models. We evaluate performance on a subset of $200$ test images across varying perturbation dimensions and magnitudes, with results shown in Figure 2. Model details and probabilistic guarantees are summarized in Table 1, where we show the following $\langle\epsilon,\ell,m\rangle$ guarantee [^3] The surrogate-based technique in hashemiscaling could not handle our perturbation dimensions.:
@@
In these experiments, we studied an $r$-dimensional darkening adversary (see Figure 1), where $r^{\prime}$ pixels in an image $x$ with intensity above $150/255$ in all channels are randomly selected for perturbation ($r=nc\times r^{\prime}$). Each direction $x_{i}^{\text{noise}}$ corresponds to darkening a channel of one such pixels, and $x^{\text{adv}}$ is parameterized by an $r$-dimensional coefficient vector $\lambda\in[\underline{\lambda},\overline{\lambda}]$. Here, $\overline{\lambda}$ induces full darkening (intensity zero), while $\underline{\lambda}$ applies partial darkening. This bound defines the perturbation space $x^{\text{adv}}\in\mathbf{I}$. We also present a demo for the status of all pixels in the segmentation mask in Figures 5,6 and 7.
@@
The degree of conservatism can be assessed by comparing $\underline{\hat{y}},\overline{\hat{y}}$] with $\underline{y},\overline{y}]$. To measure this, we compute $\mathsf{bound\_ratio}=\sum_{k=1}^{h\times w\times L}(\overline{\hat{y}}(k)-\underline{\hat{y}}(k))/\sum_{k=1}^{h\times w\times L}(\overline{y}(k)-\underline{y}(k))$ and due to the high cost of $10^{6}$ inferences on the models,we only perform the conservatism analysis on one specific case with BiSeNet from Figure 2, Table 3 in Appendix D details the numerical results.
@@ 5 Conclusion
Our verification approach is designed to be scalable and efficient. Although it does not provide deterministic guarantees, it offers strong probabilistic assurances in regions where deterministic methods are computationally intractable. This effectiveness is further demonstrated through the numerical experiments included in the paper.
Appendix A A Comparison Between application of CI and Randomized Smoothing on Robustness Analysis of SSNs.
In this section, we present a comparison with the works of fiacchini2021probabilistic; anani2024adaptive on the Cityscapes dataset cordts2016cityscapes. Since the formulation of verification guarantees differs slightly across these approaches, we first provide a general overview of the techniques proposed in fiacchini2021probabilistic; anani2024adaptive, followed by a brief description of CI-based technique. This will clarify how the comparison can be meaningfully established.
Overview 1. In fischer2021scalable, the authors introduced a probabilistic method to verify Semantic Segmentation Neural Networks. Inspired by cohen2019certified, they introduced randomness via a Gaussian noise $\nu\sim\mathcal{N}(0,\sigma^{2})\in\mathbb{R}^{h\times w\times nc}$, applied to the input, and constructed a smoothed version of the segmentation model, denoted $\overline{\mathsf{SSN}}(x)(i,j)$, for each mask pixel $(i,j)$:
@@
This leads to a unified radius $r=\sigma\Phi^{-1}(\tau)$ for the input space. While this simplifies the formulation, it introduces the concept of abstained or uncertifiable pixels—those that cannot meet the desired confidence threshold. To mitigate the issue of multiple comparisons (i.e., union bounds), the authors propose statistical corrections such as the Bonferroni and Holm–Bonferroni methods ce1936pubblicazioni; holm1979simple. Assuming the set of certifiable pixels is defined as $\mathbf{CERT}=\left\{(i,j)\mid(i,j)\text{ is certifiable}\right\}$, they propose the final guarantee in the following form:
@@
This method is referred to as $\mathsf{SEGCERTIFY}$. A key limitation of this technique is the presence of abstained pixels. The authors in anani2024adaptive addressed this by proposing an adaptive approach, $\mathsf{ADAPTIVECERTIFY}$, which reduces the number of uncertifiable pixels. However, the guarantees proposed in fiacchini2021probabilistic; anani2024adaptive apply only to the smoothed model, not the base model. To express this in terms of the base model, we define the following problem:
@@
Overview 2. In CI approach, for a given image $x$ and input set $\mathbf{B}_{r}(x)$, we perform a probabilistic reachability analysis with a $\langle\epsilon,\ell,m\rangle$ guarantee. Due to the presence of conservatism in the reachability technique, some mask pixels are marked as robust (i.e., certifiable), while others cover multiple classes and are considered unknown or uncertifiable. Let $\mathbf{CONFORMAL\_CERT}=\left\{(i,j)\mid(i,j)\text{ is certifiable}\right\}$. Then the verification objective is:
Problem 2. Given an image $x$, input set $\mathbf{B}_{r}(x)$, and sampling distribution $x^{\prime}\stackrel{{\scriptstyle\mathcal{W}}}{{\sim}}\mathbf{B}_{r}(x)$, for coverage level $\delta_{1}=1-\epsilon\in[0,1]$, hyper-parameters $\ell,m$ and confidence level $\delta_{2}=1-\mathsf{betacdf}_{\delta_{1}}(\ell,m+1-\ell)$, we aim to show the following $\langle\epsilon,\ell,m\rangle$ guarantee, where the set of certifiable pixels $\mathbf{CONFORMAL\_CERT}$ will be also determined through the verification process:
@@
Comparison. The unknown pixels in the CI-based technique are conceptually equivalent to the abstained pixels in fiacchini2021probabilistic; anani2024adaptive. Thus, given the same input set $\mathbf{B}_{r}(x)$ where $r$ is provided by fiacchini2021probabilistic, the comparison focuses to show which technique results in fewer uncertifiable mask pixels. To this end, we replicate the setup of Table 1 from anani2024adaptive, using the same HrNetV2 model wang2020deeptrained on the Cityscapes dataset with the HrNetV2-W48 backbone. In this case study, the results of the Naive approach were approximately equivalent to those of the surrogate-based approach. Therefore, we use the Naive technique for comparison, as it is more efficient than the surrogate-based method. We compare the number of uncertifiable pixels produced by the naive approach against the abstained pixels reported in their results. Our findings are summarized in Table 2, which shows the average over 200 test images. We also report the verification runtime for completeness. It is important to note that although the number of uncertifiable pixels in CI-based method is significantly smaller than in prior approaches, the CI-based formulation of guarantees also differs. In particular, we assume a prior distribution for the uncertainty, where $\mathcal{W}$ is considered to be uniform. This assumption, however, can be readily extended to worst-case distribution analysis by replacing conformal inference with robust conformal inference cauchois2020robust, a technique that strengthens guarantees at a modest cost. We leave this extension to future work.
Table 2: Percentage of uncertifiable pixels under different $(\sigma,\tau,r)$ settings for fiacchini2021probabilistic, anani2024adaptive, and the Naive approach. The verification runtime is $210$ seconds for all experiments.
$\sigma$ $\tau$ $r$ fiacchini2021probabilistic (%) anani2024adaptive (%) Naive approach (%)